Jaime age in 5 years will be the double of the age he had 6 years ago whats his age now?

Answer by Maths68(1474)

Let
Present age of Jaime = j years old

In 5 years Jaime will be = j+5 years old
6 years ago Jaime was = j-6 years old
Jaime age in 5 years will be the double of the age he had 6 years ago
j+5=2(j-6)
j+5=2j-12
j-2j=-12-5
-j=-17
Cancel out (-)
j=17


Present age of Jaime = j = 17 years old